# Integral modelling and Reinforcement Learning control of 3D liquid metal coating on a moving substrate

### v1

@@ Abstract
Metallic coatings are used to improve the durability of metal surfaces, protecting them from corrosion. These protective layers are typically deposited in a fluid state via a liquid film. Controlling instabilities in the liquid film is crucial for achieving uniform and high-quality coatings. This study explores the possibility of controlling liquid films on a moving substrate using a combination of gas jets and electromagnetic actuators. To model the 3D liquid film, we extend existing integral models to incorporate the effects of electromagnetic actuators. The control strategy was developed within a reinforcement learning framework, where the Proximal Policy Optimization (PPO) algorithm interacts with the liquid film via pneumatic and electromagnetic actuators to optimize a reward function, accounting for the amplitude of the instability waves through a trial and error process. The PPO found an optimal control law, which successfully reduced interface instabilities through a novel control mechanism, where gas jets push crests and electromagnets raise troughs using the Lorentz force.
@@ I Introduction
The interaction between liquid films and external magnetic fields is a relevant physical phenomenon for many engineering applications. Described by magnetohydrodynamic (MHD) equations, this interaction is important for the analysis of magnetic soap films [1], the evaluation of magnetic wiping solutions in dip coating processes [2, 3], the control of mould flow during continuous steel casting [4], and the design of liquid metal films for plasma confinement and heat removal inside nuclear fusion reactors [5, 6].
@@
The complexity of MHD equations poses significant challenges for numerical simulations of these flows [7], often requiring customized software implementations with prohibitive computational costs [8, 9, 10]. To avoid these limitations, Smolentsev and Abdou [11] proposed simplified MHD models based on the long-wave assumption. Other authors used models of reduced-dimensionality with a single evolution equation for film thickness based on depth-averaged equations and a gradient expansion of the velocity components truncated at first or second order [1, 8, 12].
@@
Reduced-dimensional models reproduce the relevant features of the liquid film dynamics with simplified governing equations, which enables analytical analysis and cost-effective computational simulations. Shkadov [13] and Kapitza [14] developed a 2D Integral Boundary Layer (IBL) model for moderate Reynolds number flows in terms of film thickness and flow rate. This model outperformed the single-equation model in capturing the film dynamics. Widely used in the falling liquid film community, a vast spectrum of reduced-dimensional models was developed for different Reynolds and Kapitza numbers of flows, using weighted residual methods and referred to WIBL models (reviewed by Ruyer-Quil et al. [15] and Kalliadasis et al. [16]).
The WIBL model was extended to account for the motion of the substrate and the effect of the pressure gradient and the shear stress distributions on the free surface [17]. Validated against Direct Numerical Simulations (DNS) using the Volume Of Fluid (VOF) technique to track the free surface [18], this enhanced model was used to reproduce the dynamics of the liquid film in the hot-dip galvanizing process with impinging gas jets.
@@
In this process, a metal sheet is coated with a thin layer of zinc by withdrawing it from a bath of molten zinc. The thickness of the final coating is regulated by impinging gas jets, which remove excess liquid from the surface [19, 20]. At withdrawal velocities exceeding 2 m/s, the gas jets oscillate, inducing a liquid film instability known as undulation [20, 21, 22]. This instability manifests itself as nearly 2D travelling waves [18, 23, 24], leading to an uneven coating layer that reduces the quality of the final product, limits the maximum production rate and results in substantial energy and material waste.
A feedback regulator with gas jets and electromagnets represents a possible solution to the undulation problem. Electromagnetic actuators are a new concept recently explored in developing liquid walls for nuclear fusion reactors [25, 26]. Preliminary experiments have shown that a transverse magnetic field can stabilize a liquid film of Galistan flowing over an inclined plate. In an open-loop configuration, Mirhoseini et al. [25] demonstrated the stabilizing effect of a uniform magnetic field. In a closed-loop configuration, they showed that a simple feedback system, using liquid film height observations, can locally stabilize the film by controlling the electric current in the bulk.
This work focuses on closed-loop control of coating non-uniformities in liquid metal films along a moving substrate, using gas jets and electromagnetic actuators with limited free-surface observation points. The optimal control function is developed using Proximal Policy Optimization (PPO) [27], a reinforcement learning (RL) algorithm [28]. RL methods are gaining popularity within the fluid dynamics community across a wide range of applications [29, 30, 31], particularly for high-dimensional control problems with significant noise and uncertainty, as seen in coating control for industrial galvanizing lines. A comparative analysis of RL methods and black-box optimization for flow control is presented in [32]. PPO was used in [33] to stabilize waves in a 2D falling liquid film over an inclined surface, using localized pressure pulsations and limited liquid film thickness observations without relying on knowledge of the governing equations. This work extends the 3D IBL model derived by Ivanova et al. [34], accounting for pneumatic and electromagnetic actuators. Moreover, we investigate the performances of a PPO-trained agent for wave attenuation on a liquid film over a moving substrate.
@@
The rest of the article is structured as follows. Section II outlines the problem and introduces the relevant physical quantities. Section III reviews the reference quantities used to scale the governing equations. The modelling of gas jets and electromagnetic actuators is presented in Section IV. Section V details the numerical methods for solving the reduced-dimensional model equations. Section VII presents the wave dynamics and control performance results. Finally, conclusions and perspectives are discussed in Section VIII.
@@ II Problem description and physical quantities
We consider a 3D liquid film of thickness $h$ flowing over a substrate moving against gravity $g$ at constant speed $U_{p}=2.5\;\rm{m/s}$ in contact with air, under the action of external control gas jets and electromagnetic actuators.
Figure 1: Scheme of the liquid film flowing over a substrate moving against gravity under the effect of gas jet and electromagnetic actuators.
For clarity, the physical quantities related to the liquid are written without a subscript, while those related to the air are denoted by the subscript “g”. We consider a liquid film of molten zinc with density $\rho=6570\,\rm{kg/m}^{3}$, dynamic viscosity $\mu=3.5\,\rm{mPa\,s}$, kinematic viscosity $\nu=\mu/\rho$, surface tension $\sigma=700\,\rm{mN/m}$, electrical conductivity $\sigma_{M}=2.7\,\rm{MS}/\rm{m}$ (evaluated at $450^{\circ}\,$ [35]), magnetic permeability $\mu_{\rm M}=1.26\,\mu\rm{N/A}^{2}$, magnetic diffusivity $\eta_{M}=1/(\mu_{M}\sigma_{M})=0.3\,\rm{m}^{2}/\rm{s}$ and magnetic susceptibility $\chi=(\mu_{\rm M}/\mu_{\rm M0}-1)=2.7\times 10^{-3}$ where $\mu_{M0}$ is the magnetic permeability of the void. The air has magnetic permeability $\mu_{\rm Mg}=1.26\;\mu{\rm N/A}^{2}$ and magnetic susceptibility $\chi_{\rm g}=(\mu_{\rm Mg}/\mu_{\rm M0}-1)=3.73\times 10^{-7}$.
@@
$\mu_{\rm Mg}=\frac{\chi_{\rm g}+1}{\chi+1}\mu_{\rm M}$ (1)
Figure 1 illustrates the coordinate system $\{\mathbf{O}$; $x,y,z\}$ with the $x$ axis aligned with the gravity vector $g$ pointing downward, the $y$ axis pointing along the normal direction of the wall towards the free surface, and the $z$ axis spanning the direction of flow, with origin $\mathcal{O}$ on the substrate. Due to the translational invariance of the problem, specifying a precise origin for the $x$ axis is unnecessary.
A local Cartesian coordinate system is defined on the free surface $y-h(x,z,t)=0$ tangent plane as $\{\mathbf{O}$; $\mathbf{n},\mathbf{t}_{x},\mathbf{t}_{z}\}$, with $\mathbf{n}$ representing the normal vector pointing towards the air, and $\mathbf{t}_{x}$ and $\mathbf{t}_{z}$ representing the tangential directions along $x$ and $z$ respectively. These vectors are expressed in ($\mathcal{O}$; $x,y,z$) with the following relations:
@@ III Integral Modelling with Magnetic Actuators
This section outlines the derivation of the 3D Integral Boundary Layer (IBL) model, incorporating the effects of gas jets and electromagnetic actuators. The model is presented in dimensionless form, based on the long-wave assumption [16, 15]. Subsection III.1 introduces dimensionless parameters and groups. For conciseness, the complete magnetohydrodynamic equations and boundary conditions are provided in Appendix A. Subsection III.3, introduces the boundary layer equations obtained from the long-wave approximation. Finally, Subsection III.2 discusses the steady-state solutions used to derive the closure relations for the IBL model, presented in Subsection III.4.
@@ III.1 Reference quantities and nondimensional groups
The liquid film governing equations, reported in Appendix A, are scaled as in Ivanova et al. [34], with the addition of two dimensionless numbers accounting for the electromagnetic force. In the following, all dimensionless quantities are denoted with a hat $\hat{\cdot}$. These are obtained by dividing dimensional variables by their reference quantities, identified by the subscript ‘$\rm ref$’. The streamwise and spanwise velocities are scaled by the substrate velocity $U_{p}$, and the film thickness is scaled based on the thickness arising from the viscous-gravity balance in steady-state condition [17], namely
@@
$b_{\rm ref}={\rm sup}(|b|),\qquad\qquad j_{\rm ref}=\sigma_{M}U_{p}b_{\rm ref}\,.$ (9a)
The capillary number Ca introduced with the definition of the film parameter, can be expressed as $\mbox{Ca}=\mbox{Re}^{2/3}\mbox{Ka}$, where $\mbox{Ka}=\sigma/(\rho g^{1/3}\nu^{4/3})$ is the Kapitza number and Re the Reynolds number, which is defined as:
$\mbox{Re}=\frac{\rho u_{\rm ref}h_{\rm ref}}{\mu}=\sqrt{\frac{\rho U_{p}^{3}}{g\mu}}\,.$ (10)
@@
[…]
@@ III.2 Flat-film steady state solution
Before deriving the reduced-dimensional model, it is worth analysing the steady-state flat film solution of the governing equations detailed in Appendix A, excluding the influence of external gas jet actuators. Assuming a stable ($\partial_{t}h=0$) and flat ($\partial_{x}=0\wedge\partial_{z}=0$) surface, without external shear stress ($\tau_{g,x}=\tau_{g,z}=0$) and no external pressure ($p_{g}=0$) distributions, and considering a uniform transversal magnetic field $b$, the flow field becomes parallel ($w=0$) and unidimensional ($v=0$). The Navier-Stokes equation reduces to the balance of viscous stresses, gravity, and Lorentz force along the streamwise direction $x$:
@@
The non-dimensional steady-state velocity $\bar{u}(\hat{y})$ of (13) with (14) reads:
$\bar{u}(\hat{y})=n_{1}e^{\left(\mbox{Ha}\;\hat{b}\hat{y}\right)}+n_{2}e^{\left(-\mbox{Ha}\;\hat{b}\hat{y}\right)}+\frac{1}{\mbox{Ha}^{2}\hat{b}^{2}},$ (15)
@@
$n_{1}=-\left(1+n_{2}+\frac{1}{\mbox{Ha}^{2}\hat{b}^{2}}\right),$ (16a)
$n_{2}=-\frac{e^{\left(\hat{h}\mbox{Ha}\;\hat{b}\right)}\left(1+\frac{1}{\mbox{Ha}^{2}\hat{b}^{2}}\right)}{2\cosh{\left(\mbox{Ha}\;\hat{b}\hat{h}\right)}}.$ (16b)
@@
Equations (17) and (18) describe a family of flat-film solutions that depend on the flat-film thickness $\bar{h}$. This family of solutions and their asymptotic behaviours is discussed in Section VII.1.
@@ III.3 The long wave formulation
The first step in the derivation of the reduced-dimensional model consists in scaling the magnetohydrodynamic equations and the boundary conditions presented in Appendix A using the reference quantities introduced in Subsection III.1. Retaining only terms up to $O(\varepsilon)$ and assuming the surface tension terms and Re to be of $O(1)$, results in the following boundary layer equations:
$\begin{split}\delta\big{(}\partial_{\hat{t}}\hat{u}+\hat{u}\partial_{\hat{x}}\hat{u}+\hat{v}\partial_{\hat{y}}\hat{u}&+\hat{w}\partial_{\hat{z}}\hat{u}\big{)}=-\partial_{\hat{x}}\hat{p_{x}}\\ &+\partial^{2}_{\hat{y}\hat{y}}\hat{u}+1-\mbox{Ha}^{2}\hat{b}^{2}\hat{u},\end{split}$ (19b)
@@
$\begin{split}\delta\big{(}\partial_{\hat{t}}\hat{w}+\hat{u}\partial_{\hat{x}}\hat{w}+\hat{v}\partial_{\hat{y}}\hat{w}&+\hat{w}\partial_{\hat{z}}\hat{w}\big{)}=-\partial_{\hat{z}}\hat{p_{z}}\\ &+\partial^{2}_{\hat{y}\hat{y}}\hat{w}-\mbox{Ha}^{2}\hat{b}^{2}\hat{w},\end{split}$ (19d)
with boundary conditions:
$\mathbf{\hat{v}}\big{|}_{\hat{y}=0}=(\hat{u},\hat{v},\hat{w})^{T}\big{|}_{\hat{y}=0}=(-1,0,0)^{T},$ (20a)
$\hat{v}\big{|}_{\hat{y}=\hat{h}}=\partial_{\hat{t}}\hat{h}+\hat{u}\big{|}_{\hat{y}=\hat{h}}\partial_{\hat{x}}\hat{h}+\hat{w}\big{|}_{\hat{y}=\hat{h}}\partial_{\hat{z}}\hat{h},$ (20b)
$\hat{p}\big{|}_{\hat{y}=\hat{h}}=\hat{p}_{g}-\frac{\varepsilon^{3}}{\mbox{Ca}}(\partial_{\hat{x}\hat{x}}\hat{h}+\partial_{\hat{z}\hat{z}}\hat{h})-\frac{\Lambda}{2}\hat{b},$ (20c)
$\partial_{\hat{y}}\hat{u}\big{|}_{\hat{y}=\hat{h}}=\Xi,\qquad\qquad\partial_{\hat{y}}\hat{w}\big{|}_{\hat{y}=\hat{h}}=\Psi,$ (20d)
where $\Lambda\hat{b}/2$ represent the normal component of the Maxwell stresses, with $\Psi$ and $\Xi$ accounting for the gas-jet induced shear stress $\hat{\bm{\tau}}_{g}$ and the tangential component of the Maxwell stresses $\hat{\bm{\tau}}_{M}=(\hat{b}^{2}\partial_{\hat{x}}\hat{h},\hat{b}^{2}\partial_{\hat{z}}\hat{h})$ at the free surface, reading:
$\Psi=\hat{\tau}_{g,z}+\Lambda\,\hat{b}^{2}\partial_{\hat{z}}\hat{h},\qquad\Xi=\hat{\tau}_{g,x}+\Lambda\,\hat{b}^{2}\partial_{\hat{x}}\hat{h},$ (21)
@@
The integral boundary layer model is derived by integrating the boundary layer equations (19) over the liquid film thickness ($\hat{h}$) using the Leibniz integral rule with the boundary conditions (20). This results in a set of nonlinear hyperbolic partial differential equations in terms of $\hat{h}$, $\hat{q}_{x}$, and $\hat{q}_{z}$, reading:
$\partial_{\hat{t}}\mathbf{t}+\nabla\cdot\mathbf{F}=\mathbf{s},$ (22)
where $\mathbf{t}=(\hat{h},\hat{q}_{x},\hat{q}_{z})^{\text{T}}$, $\mathbf{F}$ is the flux matrix defined as:
@@
$\mathbf{s}&=\delta^{-1}\begin{pmatrix}0\\ \hat{h}\Big{(}-\partial_{\hat{x}}\hat{p}_{g}+\partial_{\hat{x}\hat{x}\hat{x}}\hat{h}+\partial_{\hat{x}\hat{z}\hat{z}}\hat{h}-\frac{\Lambda}{2}\partial_{\hat{x}}\hat{b}\Big{)}\\ \hat{h}\Big{(}-\partial_{\hat{z}}\hat{p}_{g}+\partial_{\hat{z}\hat{z}\hat{z}}\hat{h}+\partial_{\hat{z}\hat{x}\hat{x}}\hat{h}-\frac{\Lambda}{2}\partial_{\hat{z}}\hat{b}\Big{)}\end{pmatrix}+\\ &+\delta^{-1}\begin{pmatrix}0\\ \Lambda\,\hat{b}^{2}\partial_{\hat{x}}\hat{h}+\hat{h}+\Delta\hat{\tau}_{x}-\mbox{Ha}^{2}\hat{b}^{2}\hat{q}_{x}\\ \Lambda\,\hat{b}^{2}\partial_{\hat{z}}\hat{h}+\Delta\hat{\tau}_{z}-\mbox{Ha}^{2}\hat{b}^{2}\hat{q}_{z}\end{pmatrix},\end{aligned}$
@@
The convective terms $F_{i,j}$ with $j>1$, along with the wall-shear stress component, are expressed in terms of $\hat{h}$, $\hat{q}_{x}$, and $\hat{q}_{z}$ through closure relations. These relations are based on the assumption of self-similar velocity profile in streamwise and spanwise directions [38, 39]. The shape of these profiles reflects that of the leading-order steady-state solutions given in (15), here concisely rewritten as:

### v2

@@ Abstract
Metallic coatings are used to enhance the durability of metal surfaces by protecting them from corrosion. These protective layers are typically deposited in a fluid state via a liquid film. Controlling instabilities in the liquid film is crucial to achieving uniform, high-quality coatings. This study explores the possibility of controlling liquid films on a moving substrate using a combination of gas jets and electromagnetic actuators. To model the 3D liquid film, we extend existing integral models to incorporate the effects of electromagnetic actuators. The control strategy was developed within a reinforcement learning framework, in which the Proximal Policy Optimisation (PPO) algorithm interacts with the liquid film via pneumatic and electromagnetic actuators to optimise a reward function that accounts for instability-wave amplitude through a trial-and-error process. The PPO identified an optimal control law that reduced interface instabilities via a novel mechanism: gas jets push crests, and electromagnets raise troughs via the Lorentz force.
@@ I Introduction
The interaction between liquid films and external magnetic fields is a relevant physical phenomenon with many engineering applications, including the analysis of magnetic soap films [28], the evaluation of magnetic wiping solutions in dip coating processes [13, 41], the control of mould flow during continuous steel casting [29], and the design of liquid metal films for plasma confinement and heat removal inside nuclear fusion reactors [56, 38].
@@
The complexity of the magnetohydrodynamic MHD equations, describing the liquid and electromagnetic dynamics, poses significant challenges for numerical simulations [36], often requiring customised software implementations with prohibitive computational costs [30, 50, 37]. To avoid these limitations, Smolentsev and Abdou [49] derived a simplified MHD model based on the long-wave assumption. Other authors used reduced-dimensional models with a single evolution equation for film thickness, based on depth-averaged equations and a gradient expansion of the velocity profile [28, 30, 46].
@@
Reduced-dimensional models are commonly used to capture the relevant features of liquid film dynamics by employing simplified governing equations, thereby reducing computational costs. Shkadov [48] and Kapitza [26] developed a 2D Integral Boundary Layer (IBL) model for moderate Reynolds number flows in terms of film thickness and flow rate, which outperformed the single-equation model in capturing the film dynamics.
A two-dimensional IBL model describing a liquid film flowing over a moving substrate under external pressure and shear-stress distributions at the free surface was first introduced by Mendez et al. [32]. The formulation was later extended to three dimensions to account for spanwise effects [22], and its predictions were validated against Direct Numerical Simulations (DNS) of the full governing equations [3]. This 3D IBL framework was subsequently employed to reproduce the dynamics of liquid films in the hot-dip galvanising process with impinging gas jets.
@@
The hot-dip galvanising process consists in coating a metal strip with a thin layer of zinc to protect it against corrosion. The zinc layer is applied by immersing the metal sheet in molten zinc and then withdrawing it vertically. As the sheet is withdrawn, it entrains a thin layer of liquid zinc, which then solidifies into a protective coating. The thickness of the final coating is controlled by impinging gas jets that remove the excess liquid coming from the bath [8, 19]. For withdrawal velocities above 2 m/s, the gas jets begin to oscillate, inducing an instability in the liquid film known as undulation [19, 33, 31]. This instability manifests as nearly two-dimensional travelling waves [3, 2, 1], resulting in an uneven coating layer and reduced surface quality.
Developing active control laws to suppress undulation waves is essential for achieving uniform coatings and reducing material and energy waste. Various strategies have been proposed for liquid-film systems, such as modulating electric fields [55] or controlling mass injection at the solid boundary [51]. A particularly attractive alternative is the use of blowing gas jets and electromagnetic actuators. Electromagnetic actuation, in particular, has recently shown promise for controlling liquid metal films in plasma-confinement applications. Experiments with Galinstan films flowing over inclined surfaces under transverse magnetic fields demonstrated effective stabilisation in both open- and closed-loop configurations using film-height feedback [34, 35], highlighting its potential for undulation control in coating flows.
Optimal control strategies for liquid-film flows have been explored using linear stability analysis [51], Linear Quadratic Regulator (LQR) techniques [21], optimal-control formulations [7, 52], and Model Predictive Control (MPC) methods [55]. However, applying such model-based approaches to hot-dip galvanising remains challenging due to strong noise, substantial measurement and actuation uncertainties, and unmodelled effects—such as high Reynolds numbers, thermal gradients, oxidation, and substrate vibration, that extend beyond the range of simplified models. In contrast, Reinforcement Learning (RL) [53] offers a model-free framework for developing robust feedback laws by directly interacting with the process and optimising a reward function through trial and error. RL is inherently resilient to modelling errors and uncertainties, and has shown strong performance in fluid-dynamics applications [54, 57, 44, 42], including the stabilisation of two-dimensional falling films using only local pressure actuation and limited film-thickness observations [5].
This study introduces a novel three-dimensional IBL model describing the dynamics of a liquid zinc film flowing over a vertically moving substrate under the influence of an external magnetic field and impinging gas jets. Building on the formulation of Ivanova et al. [22], the model incorporates Maxwell stresses at the free surface and Lorentz forces within the bulk. It is further employed to design optimal feedback control laws for gas jets and electromagnetic actuators to mitigate undulation instability. The resulting control strategy is derived using the Reinforcement Learning Proximal Policy Optimisation (PPO) algorithm [45].
@@
The remainder of this article is organised as follows. Section II defines the problem and introduces the relevant physical quantities. Section III outlines the reference quantities employed to nondimensionalize the governing equations. The modelling of gas jets and electromagnetic actuators is described in Section IV. Section V details the numerical methods used to solve the reduced-dimensional model equations. The results on wave dynamics and control performance are presented in Section VII. Finally, Section VIII provides concluding remarks and future perspectives.
@@ II Problem description and physical quantities
We consider a 3D liquid film of thickness $h$ flowing over a substrate moving against gravity $g$ at constant speed $U_{p}=2.5\,$m/s in contact with air, under the action of external gas jets and electromagnetic actuators.
Figure 1: Scheme of the liquid film flowing over a substrate moving against gravity under the effect of a gas jet and electromagnetic actuators.
For clarity, the physical quantities related to the liquid are written without a subscript, while those related to the air are denoted by the subscript ”g”. We consider a liquid film of molten zinc with density $\rho=6570\,\rm{kg/m}^{3}$, dynamic viscosity $\mu=3.5\,\rm{mPa\,s}$, kinematic viscosity $\nu=\mu/\rho$, surface tension $\sigma=700\,\rm{mN/m}$, electrical conductivity $\sigma_{M}=2.7\,\rm{MS}/\rm{m}$ (evaluated at $450^{\circ}\,$C [17]), magnetic permeability $\mu_{\rm M}=1.26\,\mu\rm{N/A}^{2}$, magnetic diffusivity $\eta_{M}=1/(\mu_{M}\sigma_{M})=0.3\,\rm{m}^{2}/\rm{s}$ and magnetic susceptibility $\chi=(\mu_{\rm M}/\mu_{\rm M0}-1)=2.7\times 10^{-3}$ where $\mu_{M0}$ is the magnetic permeability of the void. The air has magnetic permeability $\mu_{\rm Mg}=1.26\;\mu{\rm N/A}^{2}$ and magnetic susceptibility $\chi_{\rm g}=(\mu_{\rm Mg}/\mu_{\rm M0}-1)=3.73\times 10^{-7}$.
@@
$\mu_{\rm Mg}=\frac{\chi_{\rm g}+1}{\chi+1}\mu_{\rm M}\,.$ (1)
Figure 1 illustrates the coordinate system $\{\mathbf{O}$; x,y,z}x,y,z\} with the $x$ axis aligned with the gravity vector $g$ pointing downward, the $y$ axis pointing along the normal direction of the wall towards the free surface, and the $z$ axis spanning the direction of flow, with origin $\mathcal{O}$ on the substrate. Because of the problem’s translational invariance, specifying a precise origin for the $x$-axis is unnecessary.
A local Cartesian coordinate system is defined on the free surface yh(x,z,t)=0y-h(x,z,t)=0 tangent plane as $\{\mathbf{O}$; $\mathbf{n},\mathbf{t}_{x},\mathbf{t}_{z}\}$, with $\mathbf{n}$ representing the normal vector pointing towards the air, and $\mathbf{t}_{x}$ and $\mathbf{t}_{z}$ representing the tangential directions along $x$ and $z$ respectively. These vectors are expressed in ($\mathcal{O}$; $x,y,z$) with the following relations:
@@ III Integral Modelling with Magnetic Actuators
   $\Lambda$    $\varepsilon\,\mbox{S}\,\mbox{Ha}^{2}/\mbox{R\textsubscript{m}}$    9.3
@@
This section introduces the reference quantities and nondimensional groups (Subsection III.1), the steady-state flat film solution under a uniform magnetic field (Subsection III.2), the first-order boundary layer equations (Subsection III.3), and the derivation of the 3D IBL model (Subsection III.4).
@@ III.1 Reference quantities and nondimensional groups
The liquid film governing equations, reported in Appendix A, are scaled as in Ivanova et al. [22], with the addition of two dimensionless numbers accounting for the electromagnetic force. In the following, all dimensionless quantities are denoted with a hat $\hat{\cdot}$. These are obtained by dividing dimensional variables by their reference quantities, identified by the subscript $\rm ref$’. The streamwise and spanwise velocities are scaled by the substrate velocity $U_{p}$, and the film thickness is scaled based on the thickness arising from the viscous-gravity balance in steady-state condition [32], namely
@@
$b_{\rm ref}={\rm sup}(|b|)\qquad{\rm and}\qquad j_{\rm ref}=\sigma_{M}U_{p}b_{\rm ref}\,.$ (9a)
The capillary number Ca introduced is given by:
$\mbox{Ca}=\mbox{Re}^{2/3}\mbox{Ka},$ (10)
where Ka is the Kapitza number and Re the Reynolds number, which are defined as:
$\mbox{Ka}=\frac{\sigma}{\rho\,g^{1/3}\,\nu^{4/3}},\qquad\mbox{Re}=\frac{\rho u_{\rm ref}\,h_{\rm ref}}{\mu}=\sqrt{\frac{\rho\,U_{p}^{3}}{g\,\mu}}\,.$ (11)
@@
the magnetic Reynolds number Rm, the susceptibility ratio S, defined as:
$\mbox{R\textsubscript{m}}=\frac{u_{\rm ref}\,h_{\rm ref}}{\eta_{M}},\qquad\qquad\mbox{S}=\frac{\chi-\chi_{g}}{1+\chi_{g}},$ (13)
and the Stuart number, defined as:
$\mbox{N}=\frac{\mbox{Ha}^{2}}{\mbox{Re}}=\varepsilon\Big(\frac{t_{\rm ref}}{t_{\rm ref,M}}\Big)^{2},$ (14)
where $t_{\rm ref,M}$ is the time scale of the Loretz force, defined as:
@@
We also introduce the nondimensional parameter $\Lambda$
$\Lambda=\varepsilon\mbox{S}\mbox{Ha}^{2}/\mbox{R\textsubscript{m}},$ (16)
to characterise the intensity of the normal component of the Maxwell stresses at the free surface.
@@ III.2 Flat-film steady state solution
Before deriving the reduced-dimensional model, it is worth analysing the steady-state flat-film solution of the governing equations detailed in Appendix A, excluding the influence of external gas-jet actuators. Assuming a stable ($\partial_{t}h=0$) and flat ($\partial_{x}=0\wedge\partial_{z}=0$) surface, without external shear stress ($\tau_{g,x}=\tau_{g,z}=0$) and no external pressure ($p_{g}=0$) distributions, and considering a uniform transversal magnetic field $b$, the flow field becomes parallel ($w=0$) and unidimensional ($v=0$). The Navier-Stokes equation reduces to the balance of viscous stresses, gravity, and Lorentz force along the streamwise direction $x$:
@@
[…]
$+n_{2}\,exp(-\mbox{Ha}\;\hat{b}\;\hat{y})+\frac{1}{\mbox{Ha}^{2}\hat{b}^{2}},$
@@
$n_{1}$ $=-\left(1+n_{2}+\frac{1}{\mbox{Ha}^{2}\hat{b}^{2}}\right),$ (20)
$n_{2}$ $=-\frac{\,exp(\hat{h}\mbox{Ha}\;\hat{b})\left(1+\frac{1}{\mbox{Ha}^{2}\hat{b}^{2}}\right)}{2\cosh{\left(\mbox{Ha}\;\hat{b}\hat{h}\right)}}$.
@@
Equations (21) and (22) describe a family of flat-film solutions that depend on the flat-film thickness $\bar{h}$. This family of solutions and their asymptotic behaviours are discussed in Section VII.1.
@@ III.3 The long wave formulation
We assume that the distribution of external stresses at the free surface remains independent of the film dynamics, unless explicitly prescribed by a control law. Furthermore, advection of the magnetic field is neglected, and surface tension is assumed to be strong. These assumptions correspond to the following order-of-magnitude estimates for the magnetic Reynolds Rm and Capillary Ca numbers, reading:
$\mbox{R\textsubscript{m}}=O(\varepsilon),\qquad\qquad\mbox{Ca}=O(\varepsilon^{3}),$ (23)
with the remaining nondimensional groups assumed to be $O(1)$ quantities.
The Integral Boundary Layer (IBL) model is derived from the first-order boundary-layer equations, obtained by applying the long-wave scaling introduced in Subsection III.1 to the magnetohydrodynamic (MHD) governing equations and boundary conditions reported in Appendix A. Retaining terms up to $O(\varepsilon)$ leads to the following set of equations:
@@
where $\hat{v}=O(\varepsilon)$. The boundary conditions at the substrate $\hat{y}=0$ are given by:
$\hat{u}=-1,\qquad\qquad\hat{v}=0,\qquad\qquad\hat{w}=0,$ (25)
and at the free-surface $\hat{y}=\hat{h}$ are given by:
$\hat{v}=\partial_{\hat{t}}\hat{h}+\hat{u}\partial_{\hat{x}}\hat{h}+\hat{w}\partial_{\hat{z}}\hat{h}$, (26a)
$\hat{p}=\hat{p}_{g}-\frac{\varepsilon^{3}}{\mbox{Ca}}(\partial_{\hat{x}\hat{x}}\hat{h}+\partial_{\hat{z}\hat{z}}\hat{h})-\frac{\Lambda}{2}\hat{b}$, (26b)
$\partial_{\hat{y}}\hat{u}=\Xi,\qquad\qquad\partial_{\hat{y}}\hat{w}=\Psi,$ (26c)
[…]
$\Psi=\hat{\tau}_{g,z}+\Lambda\,\hat{b}^{2}\partial_{\hat{z}}\hat{h},\qquad\Xi=\hat{\tau}_{g,x}+\Lambda\,\hat{b}^{2}\partial_{\hat{x}}\hat{h}$. (27)
It is worth stressing that, besides the long-wave condition $\varepsilon\ll 1$, the validity of the boundary-layer approximation (and in particular the assumption $\partial_{\hat{y}}\hat{p}=0$) also requires that inertial effects remain moderate, which in turn demands the reduced Reynolds number $\delta=\varepsilon Re=O(1)$ [24]. Nevertheless, several studies have shown that integral (IBL) and weighted-integral (WIBL) models derived from (24), (25), and (26) remain accurate well beyond this formal limit, at least in the absence of magnetic effects. For instance, Denner et al. [11] validated a WIBL model up to $\delta=110.3$ ($Re=77$), Mendez et al. [32] obtained consistent results for a moving substrate at $\delta=73.9$ and $Re=478$, and Barreiro-Villaverde et al. [1] confirmed the same boundary-layer structure under jet-wiping conditions at $\delta\approx 76$. In the present configuration, the reduced Reynolds number reaches $\delta=266$, which exceeds the asymptotic range but remains within one order of magnitude of previously validated cases, supporting the practical use of the boundary-layer approximation in this study.
@@
The integral boundary layer model is derived by integrating the boundary layer equations (24) over the liquid film thickness ($\hat{h}$) using the Leibniz integral rule with the boundary conditions (25) and (26). This results in a set of nonlinear hyperbolic partial differential equations in terms of $\hat{h}$, $\hat{q}_{x}$, and $\hat{q}_{z}$, reading:
$\partial_{\hat{t}}\begin{pmatrix}\hat{h}\\ \hat{q}_{x}\\ \hat{q}_{z}\end{pmatrix}+\nabla\cdot\mathbf{F}=\mathbf{s},$ (28)
where $\mathbf{F}$ is the flux matrix defined as:
$\mathbf{F}$ $=$ $\begin{pmatrix}F_{11}&F_{21}\\ F_{12}&F_{22}\\ F_{13}&F_{23}\end{pmatrix}^{T}=\begin{pmatrix}\hat{q}_{x}&\hat{q}_{z}\\ \int_{0}^{\hat{h}}\hat{u}^{2}d\hat{y}&\int_{0}^{\hat{h}}\hat{u}\hat{w}d\hat{y}\\ \int_{0}^{\hat{h}}\hat{u}\hat{w}d\hat{y}&\int_{0}^{\hat{h}}\hat{w}^{2}d\hat{y}\end{pmatrix}^{T},$ (29)
@@
The convective terms $F_{i,j}$ with $j>1$, along with the wall-shear stress component, are expressed in terms of $\hat{h}$, $\hat{q}_{x}$, and $\hat{q}_{z}$ through closure relations. These relations are based on the assumption of a self-similar velocity profile in streamwise and spanwise directions [25, 47]. As discussed in the previous section, this assumption follows naturally from the boundary-layer structure implied by the long-wave formulation, although it formally requires moderate values of the reduced Reynolds number $\delta$. Here, we retain the traditional IBL closure approach on the same pragmatic grounds—namely, that the shape of these profiles reflects that of the leading-order steady-state solutions given in (19), here concisely rewritten as: